• Log InLog In
  • Register
Liquid`
Team Liquid Liquipedia
EDT 10:59
CEST 16:59
KST 23:59
  • Home
  • Forum
  • Calendar
  • Streams
  • Liquipedia
  • Features
  • Store
  • EPT
  • TL+
  • StarCraft 2
  • Brood War
  • Smash
  • Heroes
  • Counter-Strike
  • Overwatch
  • Liquibet
  • Fantasy StarCraft
  • TLPD
  • StarCraft 2
  • Brood War
  • Blogs
Forum Sidebar
Events/Features
News
Featured News
[ASL22] Ro4 Preview: Mirror Mirror6[ASL22] Ro8 Preview: Within Reach5[ASL22] Ro8 Preview: In A Tizzy11[ASL22] Ro16 Preview: Holy Diver5[ASL22] Ro16 Preview: Rough Waters10
Community News
Weekly Cups (October 5-11): Farewell to 5.0.162Rollback Netcode in Brood War895.0.17 Patch Notes - October 7, 202624Weekly Cups Results (Sep 28-Oct 4)0SC4ALL: II SC2 Complete Invited Player Lineup10
StarCraft 2
General
Server Blocker SCII needs to put more weight on strategy and tact HOW TO HIRE AND RECOVER LOST OR SCAMMED ETH How do you feel about the mass reverts in the 5.0.17 PTR? 5.0.17 Patch Notes - October 7, 2026
Tourneys
$3,500 WardiTV Not-Summer Championship Stellar Fest TWO the Moon (Dec 16-20) ISSL (IntoTheTV x SOOP SC2 League): Premier RSL 2026 Offline Finals LCQ Sign-Ups ScienceCraft (October 24-25) - Live Event
Strategy
[H] ZvP Mid-Late Game: Stalkers Collossi HT
Custom Maps
[G] How to playtest arcade maps [A] Nemrods 1/4 players [M] (2) Sweltering Sands [M] (2) Frigid Storage
External Content
The PondCast: SC2 News & Results Mutation # 547 Temple Of Terror Mutation # 546 Catch the Train Mutation # 545 And Drops and Rifts
Brood War
General
ASL FINALS HYPE VIDEO 2X Rollback Netcode in Brood War BW General Discussion ASL22 General Discussion BGH Auto Balance -> http://bghmmr.eu/
Tourneys
[ASL22] Semifinal B ASL 22 Finals [Megathread] Daily Proleagues [ASL22] Semifinal A
Strategy
Cliff Jump Revisited (1 in a 1000 strategy) Simple Questions, Simple Answers Replay Review Process - What do you do?
Other Games
General Games
Dawn of War IV Nintendo Switch Thread Stormgate/Frost Giant Megathread General RTS Discussion Thread Warcraft III: The Frozen Throne
Dota 2
Dota 2 Champions League Season 3 Begins April 25!
League of Legends
[TL LoL EUW IHs] Teemo shall perish
Heroes of the Storm
Heroes of the Storm 2.0
Hearthstone
Deck construction bug
TL Mafia
TL Mafia Community Thread
Community
General
US Politics Mega-thread Things Aren’t Peaceful in Palestine Russo-Ukrainian War Thread Canadian Politics Mega-thread Artificial Intelligence Thread
Fan Clubs
Serral Fan Club
Media & Entertainment
[Manga] One Piece Movie Discussion! Diablo Animated Series on Netflix
Sports
Football (Soccer) Thread TeamLiquid Health and Fitness Initiative For 2023 MLB/Baseball 2023
World Cup 2022
Tech Support
Computer Build, Upgrade & Buying Resource Thread
TL Community
Recent Gifted Posts
Blogs
The Link Between Gaming and …
TrAiDoS
[ASL22] Ro4 Day1 Ticket Giv…
bITt.mAN
38 yo Retired SWE loo…
PurE)Rabbit-SF
Can Bots Beat Pros?? Starcr…
namkraft
[meme] I finally understa…
LUCKY_NOOB
Customize Sidebar...

Website Feedback

Closed Threads



Active: 12480 users

Coins tosses - Page 5

Blogs > Sea_Food
Post a Reply
Prev 1 2 3 4 5 6 7 8 Next All
Bibdy
Profile Joined March 2010
United States3481 Posts
Last Edited: 2011-07-15 22:46:04
July 15 2011 22:37 GMT
#81
On July 16 2011 07:14 hypercube wrote:
Show nested quote +
On July 16 2011 02:38 Chill wrote:
Random:
HH x
HT o
TH o
TT x
50%

After knowing at least one was heads:
HH x
HT o
TH o
TT
66.6%

O_o

I think you're just thinking about it wrong. His knowledge isn't affecting the outcome - the outcome was random but he gave you specific information about the result that eliminates one of the possibilities.

If he flipped a coin, told you it was heads, and then flipped the other coin, there would be a 50% chance one of them was tails. But that's not the same thing. In scenario 1, TT was an option was that later eliminated with information. In scenario 2, TT is never a possibility so it doesn't factor in.


How about this:

Random:
HH -> H
HT -> T
TH -> T
TT -> H
50%

After knowing at least one was heads:
1 x HH -> H
0.5 x HT -> 0.5T
0.5 x TH -> 0.5T
TT
50%


I don't see a clear reason to prefer your distribution to mine. For example if he just looked at one of the coins in secret and told what he saw then the second distribution is the correct one (because for HT and TH he'd say tails 50% of the time).

edit: to clarify, everyone who got 2/3 was assuming that every time one of the coins comes up heads your friend will say so. If that's true you are correct to just count up the cases. However, this assumes that he has an a priori preference for heads, i.e for HT he'll always say one of them is heads and never that one of them is tails. I just don't feel like that assumption is justified, given the description of the problem.


I'm having a hard time figuring out why you're giving double the probability of HH occurring there. That's an assumption nobody made at all. The only assumption made is that your friend flipped two coins and you know that ONE of them came up heads.

The most important piece of information to realize in the entire puzzle is that your friend doesn't specify WHICH SPECIFIC coin came up Heads. Only that one of them did. Out of the four potential possibilities, the clue he gives you only rules out ONE of them, not two of them.

Hypothetically, if he told you that the FIRST coin came up heads, then your outcome list would look like this (remember that the first letter represents the first coin, and the second letter represents the second coin):

HH
HT
TH
TT

And you'd have a 50:50 chance, of picking the right answer. But, that's NOT what he told you. There's a big distinction one needs to make between him telling you that a specific coin came up heads, and that either one of them did.

This was actually part of the original question:

Your friend tosses two coins, then asks you to guess how the coins landed. You reply that you cannot know. Then your friend reveals that one of the coins he threw landed heads. Now how did the other coin land?
(Correct) answer: There is a 67% chance that the other coin landed tails. Why? Because after he told you one was heads, the remaining possibilities of the coin lands were heads-heads, heads-tails and tails-heads, therefore in two out of three cases the other coin is tails. If he told you that the first coin thrown was heads, then the chances would be 50-50, since there would be only two possibilities, heads-heads and heads-tails.
hypercube
Profile Joined April 2010
Hungary2735 Posts
Last Edited: 2011-07-15 23:12:03
July 15 2011 23:08 GMT
#82
On July 16 2011 07:37 Bibdy wrote:
Show nested quote +
On July 16 2011 07:14 hypercube wrote:
On July 16 2011 02:38 Chill wrote:
Random:
HH x
HT o
TH o
TT x
50%

After knowing at least one was heads:
HH x
HT o
TH o
TT
66.6%

O_o

I think you're just thinking about it wrong. His knowledge isn't affecting the outcome - the outcome was random but he gave you specific information about the result that eliminates one of the possibilities.

If he flipped a coin, told you it was heads, and then flipped the other coin, there would be a 50% chance one of them was tails. But that's not the same thing. In scenario 1, TT was an option was that later eliminated with information. In scenario 2, TT is never a possibility so it doesn't factor in.


How about this:

Random:
HH -> H
HT -> T
TH -> T
TT -> H
50%

After knowing at least one was heads:
1 x HH -> H
0.5 x HT -> 0.5T
0.5 x TH -> 0.5T
TT
50%


I don't see a clear reason to prefer your distribution to mine. For example if he just looked at one of the coins in secret and told what he saw then the second distribution is the correct one (because for HT and TH he'd say tails 50% of the time).

edit: to clarify, everyone who got 2/3 was assuming that every time one of the coins comes up heads your friend will say so. If that's true you are correct to just count up the cases. However, this assumes that he has an a priori preference for heads, i.e for HT he'll always say one of them is heads and never that one of them is tails. I just don't feel like that assumption is justified, given the description of the problem.


I'm having a hard time figuring out why you're giving double the probability of HH occurring there. That's an assumption nobody made at all. The only assumption made is that your friend flipped two coins and you know that ONE of them came up heads.


It's called discounting in poker. There's an equal chance of the throw coming up HH or HT. But HH leaves your friend with only one option, he can only say that the one of the coins is heads. HT leaves him with two options: he can either say that one of the coins is heads or (presumably) that one of them is tails.
The chances of the throw coming up HT and him saying one of the coins is heads is 0.25x0.5=0.125 25% for HT and 50% for choosing to point out heads instead of tails.

edit: for HH we have 0.25x1=0.25. 25% for HH and 100% for pointing out heads. 0.25 is twice as much as 0.125 hence the "double probability"

The most important piece of information to realize in the entire puzzle is that your friend doesn't specify WHICH SPECIFIC coin came up Heads. Only that one of them did. Out of the four potential possibilities, the clue he gives you only rules out ONE of them, not two of them.


It rules out one but it might also discount TH and HT by 50%, depending on our model of his behaviour.

Hypothetically, if he told you that the FIRST coin came up heads, then your outcome list would look like this (remember that the first letter represents the first coin, and the second letter represents the second coin):

HH
HT
TH
TT

And you'd have a 50:50 chance, of picking the right answer. But, that's NOT what he told you. There's a big distinction one needs to make between him telling you that a specific coin came up heads, and that either one of them did.


Yeah, that's not what I mean.

This was actually part of the original question:

"Then your friend reveals that one of the coins he threw landed heads."


That's literally all we know about our friend. He tells the truth and he has the option to tell that one of the coins is heads. Is it a stretch to assume that he could say one of them is tails if HT or TH came up? I don't think it is.
"Sending people in rockets to other planets is a waste of money better spent on sending rockets into people on this planet."
Tatum26
Profile Blog Joined July 2011
United States20 Posts
July 15 2011 23:25 GMT
#83
Coin 1 Coin 2
H1(.5) H(.5) = .25
H1(.5) T(.5) = .25
H2(.5) H(.5) = .25
H2(.5) T(.5) = .25


this is about as straight forward as it gets. H1 and H2 signify the two sides of the double-headed coin. as you can see, the probability of getting HH is 2/4, or 50%
Bibdy
Profile Joined March 2010
United States3481 Posts
July 15 2011 23:29 GMT
#84
Okay, now you've both completely lost me. I don't see what's the point in modelling the guy's behaviour when its obviously a question that's meant to challenge your ability to consider how the information presented to you only rules out one possibility, nor can I figure out where the hell a double-headed coin comes from.

I'm just going to back away, slowly.
hypercube
Profile Joined April 2010
Hungary2735 Posts
July 15 2011 23:44 GMT
#85
On July 16 2011 08:29 Bibdy wrote:
Okay, now you've both completely lost me. I don't see what's the point in modelling the guy's behaviour when its obviously a question that's meant to challenge your ability to consider how the information presented to you only rules out one possibility, nor can I figure out where the hell a double-headed coin comes from.

I'm just going to back away, slowly.


If that's what you think there's no way I can convince you.

Won't bite on the double-headed coin.
"Sending people in rockets to other planets is a waste of money better spent on sending rockets into people on this planet."
GogoKodo
Profile Blog Joined April 2003
Canada1785 Posts
Last Edited: 2011-07-16 01:59:54
July 16 2011 01:54 GMT
#86
On July 16 2011 08:08 hypercube wrote:
Show nested quote +
On July 16 2011 07:37 Bibdy wrote:
On July 16 2011 07:14 hypercube wrote:
On July 16 2011 02:38 Chill wrote:
Random:
HH x
HT o
TH o
TT x
50%

After knowing at least one was heads:
HH x
HT o
TH o
TT
66.6%

O_o

I think you're just thinking about it wrong. His knowledge isn't affecting the outcome - the outcome was random but he gave you specific information about the result that eliminates one of the possibilities.

If he flipped a coin, told you it was heads, and then flipped the other coin, there would be a 50% chance one of them was tails. But that's not the same thing. In scenario 1, TT was an option was that later eliminated with information. In scenario 2, TT is never a possibility so it doesn't factor in.


How about this:

Random:
HH -> H
HT -> T
TH -> T
TT -> H
50%

After knowing at least one was heads:
1 x HH -> H
0.5 x HT -> 0.5T
0.5 x TH -> 0.5T
TT
50%


I don't see a clear reason to prefer your distribution to mine. For example if he just looked at one of the coins in secret and told what he saw then the second distribution is the correct one (because for HT and TH he'd say tails 50% of the time).

edit: to clarify, everyone who got 2/3 was assuming that every time one of the coins comes up heads your friend will say so. If that's true you are correct to just count up the cases. However, this assumes that he has an a priori preference for heads, i.e for HT he'll always say one of them is heads and never that one of them is tails. I just don't feel like that assumption is justified, given the description of the problem.


I'm having a hard time figuring out why you're giving double the probability of HH occurring there. That's an assumption nobody made at all. The only assumption made is that your friend flipped two coins and you know that ONE of them came up heads.


It's called discounting in poker. There's an equal chance of the throw coming up HH or HT. But HH leaves your friend with only one option, he can only say that the one of the coins is heads. HT leaves him with two options: he can either say that one of the coins is heads or (presumably) that one of them is tails.
The chances of the throw coming up HT and him saying one of the coins is heads is 0.25x0.5=0.125 25% for HT and 50% for choosing to point out heads instead of tails.

edit: for HH we have 0.25x1=0.25. 25% for HH and 100% for pointing out heads. 0.25 is twice as much as 0.125 hence the "double probability"

Show nested quote +
The most important piece of information to realize in the entire puzzle is that your friend doesn't specify WHICH SPECIFIC coin came up Heads. Only that one of them did. Out of the four potential possibilities, the clue he gives you only rules out ONE of them, not two of them.


It rules out one but it might also discount TH and HT by 50%, depending on our model of his behaviour.

Show nested quote +
Hypothetically, if he told you that the FIRST coin came up heads, then your outcome list would look like this (remember that the first letter represents the first coin, and the second letter represents the second coin):

HH
HT
TH
TT

And you'd have a 50:50 chance, of picking the right answer. But, that's NOT what he told you. There's a big distinction one needs to make between him telling you that a specific coin came up heads, and that either one of them did.


Yeah, that's not what I mean.

Show nested quote +
This was actually part of the original question:

"Then your friend reveals that one of the coins he threw landed heads."


That's literally all we know about our friend. He tells the truth and he has the option to tell that one of the coins is heads. Is it a stretch to assume that he could say one of them is tails if HT or TH came up? I don't think it is.

I'm not sure where you get the option for the friend from. The problem explains a very specific scenario that has occurred and has nothing to do with the friends behavior.

Here:
I have in the real world flipped two coins while posting at Teamliquid. One of the coins has landed heads. Now, what is the chance the other coin landed tails? This is reality, these coins are sitting in front of me right now. There is no other information. How do you determine the chance in this single scenario?
twitter: @terrancem
ch33psh33p
Profile Blog Joined October 2010
7650 Posts
July 16 2011 01:57 GMT
#87
He can't be this ignorant, he has to be playing devil's advocate at this point. The question he is discussing is losing all relevance to whats ACTUALLY being discussed, and its becoming another analogy that relates to a different scenario in order to make this logic make sense.


Math trolls. =/
secret - never again
crate
Profile Blog Joined May 2009
United States2474 Posts
July 16 2011 02:05 GMT
#88
On July 16 2011 08:08 hypercube wrote:
Yeah, that's not what I mean.

Show nested quote +
This was actually part of the original question:

"Then your friend reveals that one of the coins he threw landed heads."


That's literally all we know about our friend. He tells the truth and he has the option to tell that one of the coins is heads. Is it a stretch to assume that he could say one of them is tails if HT or TH came up? I don't think it is.

Yes, it's a very big stretch. This is a math problem, not a real-life situation. The (implied) assumption in the original problem is that if at least one coin is heads, your friend will 100% of the time say that one coin is heads. Since we are considering a case where he does say one coin is heads....
We did. You did. Yes we can. No. || http://crawl.akrasiac.org/scoring/players/crate.html || twitch.tv/crate3333
RoKetha
Profile Joined July 2010
United States211 Posts
Last Edited: 2011-07-16 02:34:18
July 16 2011 02:34 GMT
#89
This is actually a problem with language and the framing of the problem. Your interpretation of the statement, "One of the coins is a head" determines the probability of the second coin being heads or tails.

If the statement being evaluated is, "Is at least one of these coins a head?" and the coin-flipper will always either answer, "One of the coins is a head," or, "It is not the case that one of the coins is a head," then the answer is that there is a 2/3 chance that one of the coins is a tail (I'd say it's misleading to say "the other" coin is a tail because a "first" or "original" coin was never specified).

If the statement the coin-flipper is answering is instead, "What side is showing on the first coin I examine?" then the answer is 50%.

For more information, see this article: http://en.wikipedia.org/wiki/Boy_or_Girl_paradox
]343[
Profile Blog Joined May 2008
United States10328 Posts
Last Edited: 2011-07-16 02:46:23
July 16 2011 02:45 GMT
#90
Hi dudes, I'd like to quote myself from earlier in the thread (in case you didn't read it?)


On July 16 2011 03:12 ]343[ wrote:
OP: unfortunately your "counterexample" is different from the original problem. Instead, it would go like this:

You have two fair coins, each with a heads and a tails side. You aren't allowed to flip 2 tails: that is, whenever you flip two tails, flip both again. The probability you will get 2 heads now is indeed 1/3.


This is what it means when we're saying "at least one of the coins turns up heads": the possibility TT is thrown out. Nothing else has changed.

EDIT: Compare to what the guy above me said: we don't pick ONE of the coins and say "oh this is heads." All we are doing is throwing out the case when both are tails.

We assume the coins are distinguishable, so H1T2 is different from H2T1. But H1H2 = H2H1... since both coins are heads.

In fact, to make it easier, we can impose an ordering on the coins being flipped. The first coin will come up H or T, the second coin will come up H or T, but not both (so no T1T2). Notice that the first coin always comes before the second coin, so there's no such thing as H2H1. So the possibilities are H1H2, H1T2, T1H2, and the probability both are heads is 1/3.
Writer
hypercube
Profile Joined April 2010
Hungary2735 Posts
July 16 2011 02:49 GMT
#91
On July 16 2011 11:05 crate wrote:
Show nested quote +
On July 16 2011 08:08 hypercube wrote:
Yeah, that's not what I mean.

This was actually part of the original question:

"Then your friend reveals that one of the coins he threw landed heads."


That's literally all we know about our friend. He tells the truth and he has the option to tell that one of the coins is heads. Is it a stretch to assume that he could say one of them is tails if HT or TH came up? I don't think it is.

Yes, it's a very big stretch. This is a math problem, not a real-life situation. The (implied) assumption in the original problem is that if at least one coin is heads, your friend will 100% of the time say that one coin is heads. Since we are considering a case where he does say one coin is heads....


Except that assumption isn't in the problem, which makes it ambiguous. It's not the only possible assumption and different assumptions will lead to different results. Read the Analysis of ambiguity section in Roketha's link, it discusses exactly this question.
"Sending people in rockets to other planets is a waste of money better spent on sending rockets into people on this planet."
hypercube
Profile Joined April 2010
Hungary2735 Posts
July 16 2011 02:57 GMT
#92
On July 16 2011 11:45 ]343[ wrote:
Hi dudes, I'd like to quote myself from earlier in the thread (in case you didn't read it?)


Show nested quote +
On July 16 2011 03:12 ]343[ wrote:
OP: unfortunately your "counterexample" is different from the original problem. Instead, it would go like this:

You have two fair coins, each with a heads and a tails side. You aren't allowed to flip 2 tails: that is, whenever you flip two tails, flip both again. The probability you will get 2 heads now is indeed 1/3.


This is what it means when we're saying "at least one of the coins turns up heads": the possibility TT is thrown out. Nothing else has changed.


That's a huge assumption. What you're doing is that you are rewriting the problem, introducing new information and then go on to prove that the answer is 2/3. It is in your problem. But it's a different problem, so it doesn't help us.
"Sending people in rockets to other planets is a waste of money better spent on sending rockets into people on this planet."
]343[
Profile Blog Joined May 2008
United States10328 Posts
July 16 2011 03:28 GMT
#93
On July 16 2011 11:57 hypercube wrote:
Show nested quote +
On July 16 2011 11:45 ]343[ wrote:
Hi dudes, I'd like to quote myself from earlier in the thread (in case you didn't read it?)


On July 16 2011 03:12 ]343[ wrote:
OP: unfortunately your "counterexample" is different from the original problem. Instead, it would go like this:

You have two fair coins, each with a heads and a tails side. You aren't allowed to flip 2 tails: that is, whenever you flip two tails, flip both again. The probability you will get 2 heads now is indeed 1/3.


This is what it means when we're saying "at least one of the coins turns up heads": the possibility TT is thrown out. Nothing else has changed.


That's a huge assumption. What you're doing is that you are rewriting the problem, introducing new information and then go on to prove that the answer is 2/3. It is in your problem. But it's a different problem, so it doesn't help us.


I'm pretty sure "at least one of the coins turns up heads" is mathematically precise. It means that both are not tails.
Writer
Chill
Profile Blog Joined January 2005
Calgary26004 Posts
July 16 2011 05:01 GMT
#94
lol this thread is a fucking train wreck. Jesus christ.
Moderator
hypercube
Profile Joined April 2010
Hungary2735 Posts
July 16 2011 05:25 GMT
#95
On July 16 2011 12:28 ]343[ wrote:
Show nested quote +
On July 16 2011 11:57 hypercube wrote:
On July 16 2011 11:45 ]343[ wrote:
Hi dudes, I'd like to quote myself from earlier in the thread (in case you didn't read it?)


On July 16 2011 03:12 ]343[ wrote:
OP: unfortunately your "counterexample" is different from the original problem. Instead, it would go like this:

You have two fair coins, each with a heads and a tails side. You aren't allowed to flip 2 tails: that is, whenever you flip two tails, flip both again. The probability you will get 2 heads now is indeed 1/3.


This is what it means when we're saying "at least one of the coins turns up heads": the possibility TT is thrown out. Nothing else has changed.


That's a huge assumption. What you're doing is that you are rewriting the problem, introducing new information and then go on to prove that the answer is 2/3. It is in your problem. But it's a different problem, so it doesn't help us.


I'm pretty sure "at least one of the coins turns up heads" is mathematically precise. It means that both are not tails.


It's also not from the OP.

"Then your friend reveals that one of the coins he threw landed heads."

This is the problem statement. The interpretation that he could have said "one of the coins landed tails" if he got TH or HT is reasonable. Others agree with me, read the boy or girl paradox article on wikipedia if you don't believe me.
"Sending people in rockets to other planets is a waste of money better spent on sending rockets into people on this planet."
HuKPOWA
Profile Blog Joined February 2011
United States1604 Posts
July 16 2011 05:42 GMT
#96
I feel bad for Chill....he is trying so hard to explain it to him lol!!!!!
waxypants
Profile Blog Joined September 2009
United States479 Posts
Last Edited: 2011-07-16 06:18:40
July 16 2011 06:15 GMT
#97
On July 16 2011 14:25 hypercube wrote:
Show nested quote +
On July 16 2011 12:28 ]343[ wrote:
On July 16 2011 11:57 hypercube wrote:
On July 16 2011 11:45 ]343[ wrote:
Hi dudes, I'd like to quote myself from earlier in the thread (in case you didn't read it?)


On July 16 2011 03:12 ]343[ wrote:
OP: unfortunately your "counterexample" is different from the original problem. Instead, it would go like this:

You have two fair coins, each with a heads and a tails side. You aren't allowed to flip 2 tails: that is, whenever you flip two tails, flip both again. The probability you will get 2 heads now is indeed 1/3.


This is what it means when we're saying "at least one of the coins turns up heads": the possibility TT is thrown out. Nothing else has changed.


That's a huge assumption. What you're doing is that you are rewriting the problem, introducing new information and then go on to prove that the answer is 2/3. It is in your problem. But it's a different problem, so it doesn't help us.


I'm pretty sure "at least one of the coins turns up heads" is mathematically precise. It means that both are not tails.


It's also not from the OP.

"Then your friend reveals that one of the coins he threw landed heads."

This is the problem statement. The interpretation that he could have said "one of the coins landed tails" if he got TH or HT is reasonable. Others agree with me, read the boy or girl paradox article on wikipedia if you don't believe me.



Thanks for that, I knew something wasn't sitting right with me.
OneOther
Profile Blog Joined August 2004
United States10774 Posts
July 16 2011 06:28 GMT
#98
Essentially the identical concept of probability:

http://www.teamliquid.net/forum/viewmessage.php?topic_id=232017
lolsixtynine
Profile Blog Joined January 2011
United States600 Posts
July 16 2011 06:39 GMT
#99
If my friend tells you that you don't know the laws of probability, what's the chance that you'll still be beating my GSTL fantasy team by 25 points?
blankspace
Profile Blog Joined June 2010
United States292 Posts
July 16 2011 06:51 GMT
#100
Ok I didn't read the whole thread but I take it that your main objection is the "unintuitive" nature of counting

HT
TH

as separate cases?

Well think of it this way: if you flip two coins the chances of getting exactly one head is twice that of getting two heads (if you don't buy this I suggest you try flipping some coins lol)

Ok now we only have two cases:
1: One head one tail (probability 2/3)
2: Two heads (probability 1/3)

(why? because the probability of case 1 must be twice that of case 2, and the sum must equal 1. Or 2x+x = 1 and solve for x)

Now you see that case one has probability 2/3

Or maybe another way, suppose I color one coin blue and one coin red (same as imposing an "ordering"). Your friend tells you that at least one coin has heads. Well then I think you would agree that we should consider two cases: blue head red tails, blue tails red heads. Now if the friend could only say whether the blue coin was heads, well then he could only speak half the time (just like monty hall problem).
Hello friends
Prev 1 2 3 4 5 6 7 8 Next All
Please log in or register to reply.
Live Events Refresh
Kung Fu Cup
20:00
2026 Week 27
WardiTV989
IntoTheiNu 905
TKL 349
LiquipediaDiscussion
[ Submit Event ]
Live Streams
Refresh
StarCraft 2
Lowko434
TKL 349
IndyStarCraft 184
ProTech152
StarCraft: Brood War
Britney 25544
Mini 1161
EffOrt 930
Horang2 912
BeSt 671
Soma 498
Larva 488
FanTaSy 435
Light 311
actioN 295
[ Show more ]
Snow 267
ZerO 247
Zeus 112
PianO 98
hero 93
Sharp 49
JYJ 47
Hyun 47
sSak 43
ToSsGirL 38
Shinee 37
Mong 33
sorry 31
Bishop 19
scan(afreeca) 18
Rock 13
Terrorterran 12
Noble 8
Dota 2
Gorgc5176
qojqva1651
Dendi1199
XaKoH 354
syndereN329
Counter-Strike
x6flipin227
Other Games
fl0m2581
singsing2309
Liquid`RaSZi942
hiko938
B2W.Neo818
byalli794
ceh9382
crisheroes255
Liquid`VortiX219
Pyrionflax164
Fuzer 55
ZerO(Twitch)13
edward11
fpsfer 2
Organizations
StarCraft: Brood War
lovetv 9
[ Show 11 non-featured ]
StarCraft 2
• AfreecaTV YouTube
• intothetv
• Kozan
• IndyKCrew
• Migwel
StarCraft: Brood War
• HerbMon 27
• BSLYoutube
• STPLYoutube
• ZZZeroYoutube
League of Legends
• Nemesis3604
Other Games
• WagamamaTV158
Upcoming Events
The PondCast
19h 1m
WardiTV Qualifier
20h 1m
Online Event
1d 20h
Korean StarCraft League
2 days
Afreeca Starleague
2 days
Rush vs Soulkey
CranKy Ducklings
2 days
BSL Open Qualifier
3 days
Sparkling Tuna Cup
3 days
IPSL
4 days
BSL Open Qualifier
4 days
[ Show More ]
WardiTV Weekly
4 days
IntoTheTV X SOOP
5 days
PiGosaur Cup
6 days
The PondCast
6 days
Kung Fu Cup
6 days
Liquipedia Results

Completed

Proleague 2026-10-13
Blizzard Classic Cup 2026
Copium Cup

Ongoing

ASL Season 22
Super Anchor Qualifying S3
Acropolis #5
HCC Season 3
Thunderpick World Champ. '26
ESL Pro League Season 24
Stake Ranked Episode 4
1win Private Club #1
Logitech G Play Connect 2026
SL StarSeries Fall 2026
FISSURE Playground #3
BLAST Open Fall 2026
Esports World Cup 2026
BLAST Bounty Summer 2026

Upcoming

Acropolis #5 - GSC
BSL Season 23
SC4ALL II: Brood War
BSL 23: Non-Korean Championship
HSC XXX
Stellar Fest 2: Lunar Cup
SC4ALL II: StarCraft II
Kung Fu Cup 2026 Grand Finals
RSL Offline Finals
eXTREMESLAND 2026
PGL Major Singapore 2026
Stake Ranked Episode 6
BLAST Rivals Fall 2026
IEM Beijing 2026
Stake Ranked Episode 5
PGL Masters Bucharest 2026
1win Private Club #2
TLPD

1. ByuN
2. TY
3. Dark
4. Solar
5. Stats
6. Nerchio
7. sOs
8. soO
9. INnoVation
10. Elazer
1. Rain
2. Flash
3. EffOrt
4. Last
5. Bisu
6. Soulkey
7. Mini
8. Sharp
Sidebar Settings...

Advertising | Privacy Policy | Terms Of Use | Contact Us

Original banner artwork: Jim Warren
The contents of this webpage are copyright © 2026 TLnet. All Rights Reserved.